Acid solution A is composed of 12.5% hydrochloric acid and acid solution B is composed of 34% hydrochloric acid. How much of each solution is needed to obtain 86 ounces of a solution that contains 25% of hydrochloric acid?
Answer by stanbon(75887)

Acid solution A is composed of 12.5% hydrochloric acid and acid solution B is composed of 34% hydrochloric acid. How much of each solution is needed to obtain 86 ounces of a solution that contains 25% of hydrochloric acid?
----------------------
Equation:
acid + acid = acid
0.125x + 0.34(86-x) = 0.25*86
---------------------------------------
Multiply thru by 1000 to get:
125x + 340*86 - 340x = 250*86
-------------------------
-115x = -90*86
------
x = 67.3 ounces (amt of 12.5% solution needed)
86-x = 18.7 ounces (amt of 34% solution needed)
-------------
Cheers,
Stan H.
-------------